Crisle:

if ab=3x+2, bc=8x-10, ac=69, and x>?

surjithayer:

what you want to do?

CripQUEZZ:

Show a ss please

surjithayer:

ab=3x+2 abc=c(3x+2)...(1) bc=8x-10 abc=a(8x-10) ...(2) from (1) and (2) a(8x-10)=c(3x+2) ac=69 \[a=\frac{ 69 }{ c }\] \[\frac{ 69 }{ c }(8x-10)=c(3x+2)\] \[\frac{ 69(8x-10) }{ 3x+2 }=c^2\] \[c^2\ge 0~(always)\] \[Hence~\frac{ 69(8x-10) }{ 3x+2 }\ge 0\] or \[\frac{ 8x-10 }{ 3x+2 }\ge 0\] so numerator and denominator are of same sign. \[let~8x-10 \ge 0\] or \[x \ge \frac{10}{8}\] \[x\ge \frac{5}{4}\] and \[3x+2 \ge 0\] \[x\ge \frac{-2}{3}\] combining \[x \ge \frac{5}{4}\] similarly you can solve for 8x-10<0 and 3x+2<0

surjithayer:

8x-10<0 \[x <\frac{ 10 }{ 8 }\] or \[x<\frac{ 5 }{4 }\] and 3x+2<0 \[x<\frac{ -2 }{3 }\] combining \[x<-\frac{ 2 }{3 }\] combinig \[x<-\frac{ 2 }{ 3 } ~or~x \ge \frac{ 5 }{ 4 }\]